OpenStudy (anonymous):

A wheel rolls uniformly without slipping on a horizontal ground.A mud particle from its topmost point gets detached from the wheel and the center of the wheel reaches above it after 1 sec.Then the radius of the wheel is?

OpenStudy (anonymous):

This is how I proceeded: Let the radius of the disc be R.   The time of flight of mud particle is 2R = (1/2)*g(t^2) => 2sqrt(R/10) = t . Let w be the angular velocity. The horizontal range of mud particle = t*2Rw. This must be equal to the velocity of Center of disc*1sec. So I get R= 0.62m. But the options are 2m , 2.5m, 4m, 4.5m. Can someone suggest me where I have went wrong.

OpenStudy (anonymous):

@experimentX

OpenStudy (experimentx):

didn't understand the last part of Q "the center of the wheel reaches above it after 1 sec."

OpenStudy (experimentx):

those two points coincide? (mud and the center of wheel)?

OpenStudy (anonymous):

the wheel runs over the mud particle after 1 second.

OpenStudy (experimentx):

let \( \omega \) be the angular velocity of wheel, then \( \omega r \) is the velocity of the wheel. the velocity of mud is \( 2 \omega r\), there is no way they make contact on air.

OpenStudy (experimentx):

the distance traveled by the mud is twice the distance traveled by wheel ... so mud particle should have fallen before 1s

OpenStudy (experimentx):

the flight time of the mud particle is half second.

OpenStudy (experimentx):

h=2r g=9.8 t=0.5 u=0

OpenStudy (anonymous):

Then r = 9.8/16 = 0.6125 m No such option is in the answer.I am also getting the same.

OpenStudy (experimentx):

what are your options?